A famous shop blends yogurt that is 3% fat with its no-fat yogurt to obtain low-fat yogurt that is 1% fat. How many pounds of regular yogurt and how many pounds of no-fat yogurt should be mixed to obtain 60 pounds of low fat yogurt?
Answer by stanbon(75887)

A famous shop blends yogurt that is 3% fat with its no-fat yogurt to obtain low-fat yogurt that is 1% fat. How many pounds of regular yogurt and how many pounds of no-fat yogurt should be mixed to obtain 60 pounds of low fat yogurt?
------------------------
Fat Equation:
fat + Fat = Fat
0.03x + 0*(60-x) = 0.01*60
Multiply thru by 100 to get:
3x + 0 = 60
x = 20 lbs (amt. of 3% yogurt in the mixture
60-x = 40 lbs (amt of 0% yogurt in the mixture)
===========================
Cheers,
Stan H.